OpenStudy (anonymous):

Evaluate this exponential expression. (0.125)^2/3

OpenStudy (anonymous):

\[(0.5)^{3*\frac{2}{3}} = (0.5)^2 = 0.25\]

OpenStudy (anonymous):

Thank you so much.

OpenStudy (anonymous):

close your question if it is solved.

OpenStudy (anonymous):

I will. When I figure out if it was solved, thank you very much.